# TriDF: Triplane-Accelerated Density Fields for Few-Shot Remote Sensing Novel View Synthesis

### v1

@@ Abstract
Remote sensing novel view synthesis (NVS) offers significant potential for 3D interpretation of remote sensing scenes, with important applications in urban planning and environmental monitoring. However, remote sensing scenes frequently lack sufficient multi-view images due to acquisition constraints. While existing NVS methods tend to overfit when processing limited input views, advanced few-shot NVS methods are computationally intensive and perform sub-optimally in remote sensing scenes. This paper presents TriDF, an efficient hybrid 3D representation for fast remote sensing NVS from as few as 3 input views. Our approach decouples color and volume density information, modeling them independently to reduce the computational burden on implicit radiance fields and accelerate reconstruction. We explore the potential of the triplane representation in few-shot NVS tasks by mapping high-frequency color information onto this compact structure, and the direct optimization of feature planes significantly speeds up convergence. Volume density is modeled as continuous density fields, incorporating reference features from neighboring views through image-based rendering to compensate for limited input data. Additionally, we introduce depth-guided optimization based on point clouds, which effectively mitigates the overfitting problem in few-shot NVS. Comprehensive experiments across multiple remote sensing scenes demonstrate that our hybrid representation achieves a 30$\times$ speed increase compared to NeRF-based methods, while simultaneously improving rendering quality metrics over advanced few-shot methods (7.4% increase in PSNR, 12.2% in SSIM, and 18.7% in LPIPS). The code is publicly available at https://github.com/kanehub/TriDF
Index Terms:
Remote sensing, novel view synthesis, neural radiance field, few-shot, triplane
@@ I Introduction
Novel view synthesis has emerged as a crucial area in computer vision, enabling enhanced 3D understanding through scene reconstruction from multi-view posed images. With the continuous development of remote sensing technology, the demand for 3D interpretation has grown significantly [1, 2, 3, 4, 5, 6, 7]. Novel view synthesis can accurately capture 3D information, including terrain characteristics and building details, which plays a vital role in 3D reconstruction, urban planning, and environmental monitoring [8, 9, 10, 11, 12].
@@
Recently, learning-based reconstruction methods [22, 23, 24, 25] have emerged as a promising alternative, driven by advances in deep learning. Neural Radiance Fields (NeRF) [26] represents a particularly groundbreaking development in this field [27]. By modeling scenes as continuous implicit functions, NeRF enables the generation of highly realistic novel view images. This breakthrough has inspired numerous NeRF-based approaches, leading to significant advances in various applications, including large-scale reconstruction [28, 29] and real-time rendering [30, 31, 32].
Most NVS methods focus on small-scale scenes, individual objects, or synthetic virtual scenes. To successfully reconstruct 3D scenes and render high-quality novel view images, they typically require dense camera arrays to capture hundreds of multi-view images as model inputs. They also rely on long iterative optimizations, leveraging cross-view supervision to ensure spatial consistency. However, for remote sensing scenes, these methods face two critical challenges. The first challenge is data acquisition: Remote sensing images are usually captured by satellites or drones. Due to limitations in flight paths or fuel [9, 33], it is difficult to obtain hundreds of densely captured multi-view images. As a result, novel view synthesis for remote sensing scenes often suffers from sparse input views, characterized by a limited number of viewpoints and low overlap between views. This sparsity can lead to overfitting on the input images. The second challenge is reconstruction efficiency: The optimization process required for novel view synthesis is time-consuming, which takes dozens of hours to reconstruct a single scene. This makes it unsuitable for real-time applications. To address these challenges, we explore efficient few-shot novel view synthesis in remote sensing scenes, with only 3 input views.
@@
In this paper, we propose TriDF, a hybrid 3D representation that achieves both fast training and high-fidelity renderings from sparse views in remote sensing scenes. Specifically, we explore the potential of the triplane representation in few-shot rendering by discretizing the scene space into three mutually orthogonal feature planes. High-frequency color features are mapped onto this compact structure, accelerating model convergence through direct optimization of explicit feature planes. Scene geometry is modeled as continuous density fields, while reference features from neighboring views are aggregated using an image-based rendering (IBR) framework to compensate for limited input information. Unlike traditional NeRF methods, we separate the modeling of volume density to accelerate convergence, taking advantage of its spatial continuity and simpler distribution patterns. Additionally, we introduce the depth-guided optimization based on 3D point clouds to address overfitting issues in few-shot rendering. Coarse 3D point clouds are generated through Structure-from-Motion (SfM) and Patchmatch stereo techniques, enabling sparse depth supervision to guide accurate scene geometry reconstruction. This approach achieves stable optimization with minimal computational overhead. We validate TriDF through extensive experiments on the LEVIR-NVS dataset [9]. As shown in Fig. 1(b), results demonstrate that our approach significantly enhances few-shot rendering quality across diverse scene types while maintaining a substantial advantage in reconstruction speed.
@@
  • We propose to model color and volume density independently. We directly optimize the explicit tri-plane with high-frequency color information and independently model volume density to reduce the computational burden of density fields, which achieves fast and efficient reconstruction.
  • Extensive experiments on remote sensing scenes demonstrate TriDF’s superior performance over various 3D representations and advanced few-shot NeRF methods in rendering quality. Moreover, TriDF achieves remarkable training efficiency, completing scene reconstruction in just 5 minutes (30$\times$ faster than NeRF-based methods) on a single consumer GPU.
@@ II-A Neural Radiance Fields
Neural rendering has advanced significantly in the field of novel view synthesis [25, 41, 42, 43, 44], with Neural Radiance Fields (NeRF) [26] emerging as the notable approach. NeRF achieves high-fidelity novel view synthesis by mapping coordinates and view directions to colors and volume densities via implicit neural representations. NeRF overcomes the resolution limitations inherent in traditional voxel-based methods by encoding continuous volumetric representations within neural networks. Inspired by NeRF, subsequent research has enhanced both the quality and efficiency of novel view synthesis while broadening its applications. NeRF++ [45] decomposes the scene space and applies inverted sphere parameterization, extending Neural Radiance Fields to 360° unbounded scenes. Mip-NeRF [46] improves anti-aliasing performance under multi-scale inputs by replacing ray sampling with cone sampling and introducing integrated positional encoding with adaptive frequency adjustment. Instant-NGP [30] achieves significant acceleration while maintaining reconstruction accuracy through the hash encoding of multi-resolution cascades and optimized ray-marching techniques. BungeeNeRF [28] enables accurate multi-scale reconstruction of urban scenes through a progressive learning model with multi-level supervision.
@@ II-B Neural Scene Representations
Other methods focus on utilizing explicit or hybrid scene representations for acceleration. Some researchers employ sparse grids for fast inference [49, 53, 54]. Plenoxels [31] extracts a NeRF variant into a sparse voxel grid with spherical harmonics, which enables significantly faster training through direct optimization of explicit representations without neural networks. TensoRF [55] models the radiance field as a 4D tensor and factorizes it into low-rank vectors and matrices, reducing both memory usage and reconstruction time. MINE [56] and ImMPI [9] integrate NeRF with Multiplane Images (MPI). Given a single input image, MINE predicts color and volume density at arbitrary depths and renders novel views using homography warping, enabling efficient reconstruction. Due to the inherent sparsity of 3D scenes, volume representations often suffer from redundancy. To address this, ED3G [57] proposes a compact triplane representation that stores features on axis-aligned planes and employs a lightweight decoder for efficient 3D generation. Similarly, TriMipRF [32] constructs a triplane representation parameterized by three mipmaps, substantially improving training speed while maintaining anti-aliasing effects. 3DGS [13], a recently developed reconstruction method, achieves high-quality real-time rendering through direct optimization of 3D Gaussian primitives. However, these methods still rely on dense input images for reconstruction and cannot achieve few-shot rendering.
II-C Few-shot NeRF
@@
Alternative approaches focus on incorporating regularization constraints [38, 59, 60] or additional supervision [40, 61, 62]. RegNeRF [37] introduces local depth smoothness constraints as regularization to enhance scene structure. DietNeRF [63] employs a pre-trained Vision Transformer (ViT) to enforce semantic consistency across views. FreeNeRF [39] implements frequency regularization to anneal positional encoding, enabling few-shot novel view synthesis with minimal computational overhead. DS-NeRF [40] utilizes sparse point clouds for depth supervision to ensure correct scene geometries. However, methods based on regularization face challenges in balancing quality and efficiency, and some methods require specific designs. In contrast, our proposed method achieves efficient, high-quality synthesis from sparse views without these limitations.
@@
Figure 2: Overview of the proposed TriDF. We introduce an efficient hybrid representation for few-shot novel view synthesis, which takes in sparse posed images and performs fast reconstruction. TriDF consists of a Tri-Plane branch and a Density Fields branch and separately predicts color and volume density. We also integrate the image-based rendering framework and depth-guided optimization based on 3D point clouds for such hybrid representations to stabilize the training process of few-shot NVS tasks.
@@ III-A Overview
Given a set of sparse multi-view remote sensing images with their corresponding camera poses, we aim to efficiently reconstruct the radiance field for rendering images from arbitrary novel views. An overview of our proposed method is shown in Fig. 2. To achieve efficient scene reconstruction, we propose a hybrid 3D representation that takes advantage of explicit triplane and implicit density fields, which separately predict the color and volume density of sampling points. High-frequency color features are mapped onto the triplane for direct optimization, and volume density is modeled as continuous density fields through MLPs, which significantly improves training efficiency. Although this hybrid representation accelerates training, its performance drops considerably when input views are sparse. To address this limitation, we incorporate the density fields into an image-based rendering framework to aggregate useful information. Furthermore, we implement a depth-guided optimization approach leveraging point clouds to improve the model’s robustness against sparse inputs while ensuring computational efficiency throughout the training process.
@@
A significant limitation of neural radiance fields with sparse input is their tendency to overfit the training views and fail to reconstruct accurate geometries, resulting in blurry synthesized images. To address this limitation, we propose a point cloud-based supervision. The approach begins by generating point clouds through Structure-from-Motion (SfM) and Patch Match Stereo (PMS) [16, 65]. Through re-projection, we extract key points and their associated depth information from the generated point clouds. These key points function as spatial anchors during the training process, providing additional geometric constraints. This supervision effectively stabilizes the optimization process since the incorporation of depth priors helps maintain geometric consistency across different viewpoints.
@@ III-B Triplane-Accelerated Density Fields
The triplane is an efficient 3D scene representation. It obtains corresponding feature vectors by projecting 3D points onto orthogonal feature planes and retrieves the color and volume density through MLP after feature aggregation. A key advantage of this representation is that it enables direct optimization of explicit triplane features, which allows for the use of lightweight MLP decoders that significantly accelerate reconstruction. Our research aims to extend the triplane representation to novel view synthesis tasks with sparse input while addressing two primary challenges: 1) Inaccurate modeling of volume density: The discrete nature of triplane representation makes it challenging to model fine scene geometry, particularly when limited supervision is provided. 2) Integration difficulties with the image-based rendering framework: The tri-plane representation struggles to effectively leverage this prior, leading to overly smoothed synthesis results.
@@ III-B1 Triplane Representation
To process these features, we employ a compact MLP that combines the triplane features ${f}_{\mathrm{tri}}$ and density features ${f}_{\mathrm{m}}$ to generate intermediate features. Subsequently, we encode view directions on the spherical harmonics basis. The encoded directions $\mathrm{SH}(\mathbf{d})$ and intermediate features ${f}_{\mathrm{base}}$ are then processed through the Color MLP to predict the colors. This pipeline can be formally expressed as:
@@ III-B2 Density Fields
where $\gamma$ represents positional encoding and ${f}_{\mathrm{m}}$ denotes density features. Due to spatial continuity, the volume density distribution along a ray is relatively simple compared to color, which often contains complex high-frequency components such as intricate textures. Therefore, modeling volume density alone is more efficient than simultaneously predicting both color and density, which leads to faster convergence.
@@
$\mathcal{L}_{\mathrm{color}}=\mathbb{E}_{\mathbf{r}}\|\hat{{C}}(\mathbf{r})-{C}(\mathbf{r})\|_{2}^{2},$ (8)
where $R_{i}$ indicates the set of sampled rays.

### v2

@@ Abstract
[…]
@@ I Introduction
Novel view synthesis has emerged as a crucial area in computer vision, enabling enhanced 3D understanding through scene reconstruction from multi-view posed images. With the continuous development of remote sensing technology, the demand for 3D interpretation has grown significantly [74, 50, 10, 35, 13, 34]. Novel view synthesis can accurately capture 3D information, including terrain characteristics and building details, which plays a vital role in 3D reconstruction, urban planning, and environmental monitoring [16, 65, 25, 51, 36, 9, 15].
@@
Recently, learning-based reconstruction methods [73, 6, 41, 70] have emerged as a promising alternative, driven by advances in deep learning. Neural Radiance Fields (NeRF) [42] represents a particularly groundbreaking development in this field [61]. By modeling scenes as continuous implicit functions, NeRF enables the generation of highly realistic novel view images. This breakthrough has inspired numerous NeRF-based approaches, leading to significant advances in various applications, including large-scale reconstruction [66, 60] and real-time rendering [43, 19].
Most NVS methods focus on small-scale scenes, individual objects, or synthetic virtual scenes. To successfully reconstruct 3D scenes and render high-quality novel view images, they typically require dense camera arrays to capture hundreds of multi-view images as model inputs. They also rely on long iterative optimizations, leveraging cross-view supervision to ensure spatial consistency. However, for remote sensing scenes, these methods face two critical challenges. The first challenge is data acquisition: Remote sensing images are usually captured by satellites or drones. Due to limitations in flight paths or fuel [65], it is difficult to obtain hundreds of densely captured multi-view images. As a result, novel view synthesis for remote sensing scenes often suffers from sparse input views, characterized by a limited number of viewpoints and low overlap between views. This sparsity can lead to overfitting on the input images. The second challenge is reconstruction efficiency: The optimization process required for novel view synthesis is time-consuming, which takes dozens of hours to reconstruct a single scene. This makes it unsuitable for real-time applications. To address these challenges, we explore efficient few-shot novel view synthesis in remote sensing scenes, with only 3 input views.
@@
In this paper, we propose TriDF, a hybrid 3D representation that achieves both fast training and high-fidelity renderings from sparse views in remote sensing scenes. Our motivation stems from the limitations of triplane representations: when supervision is insufficient, triplanes struggle to reconstruct fine-grained geometric structures. Moreover, jointly modeling color and volume density forces the neural radiance fields to learn low-frequency geometry and high-frequency appearance simultaneously, leading to competition for network capacity. This misalignment often results in suboptimal performance and slower convergence. To address these issues, we propose a hybrid explicit–implicit representation that models color and volume density independently. Specifically, we explore the potential of the triplane representation in few-shot rendering by discretizing the scene space into three mutually orthogonal feature planes. High-frequency color features are mapped onto this compact structure, accelerating model convergence through direct optimization of explicit feature planes. Scene geometry is modeled as continuous density fields, while reference features from neighboring views are aggregated using an image-based rendering (IBR) framework to compensate for limited input information. Since volume density exhibits simpler distribution patterns, this design reduces the optimization burden of the density fields and accelerates the overall training process.
Additionally, we introduce the depth-guided optimization based on 3D point clouds to address overfitting issues in few-shot rendering. Coarse 3D point clouds are generated through Structure-from-Motion (SfM) and Patchmatch stereo techniques, enabling sparse depth supervision to guide accurate scene geometry reconstruction. This approach achieves stable optimization with minimal computational overhead. We validate TriDF through extensive experiments on the LEVIR-NVS dataset [65]. As shown in Fig. 1(b), results demonstrate that our approach significantly enhances few-shot rendering quality across diverse scene types while maintaining a substantial advantage in reconstruction speed.
@@
  • We propose to model color and volume density independently. We directly optimize the explicit triplane with high-frequency color information and independently model volume density to reduce the computational burden of density fields, which achieves fast and efficient reconstruction.
[…]
@@ II-A Neural Radiance Fields
Neural rendering has advanced significantly in the field of novel view synthesis [70, 14, 56, 45, 12], with Neural Radiance Fields (NeRF) [42] emerging as the notable approach. NeRF achieves high-fidelity novel view synthesis by mapping coordinates and view directions to colors and volume densities via implicit neural representations. NeRF overcomes the resolution limitations inherent in traditional voxel-based methods by encoding continuous volumetric representations within neural networks. Inspired by NeRF, subsequent research has enhanced both the quality and efficiency of novel view synthesis while broadening its applications. NeRF++ [76] decomposes the scene space and applies inverted sphere parameterization, extending Neural Radiance Fields to 360° unbounded scenes. Mip-NeRF [2] improves anti-aliasing performance under multi-scale inputs by replacing ray sampling with cone sampling and introducing integrated positional encoding with adaptive frequency adjustment. Instant-NGP [43] achieves significant acceleration while maintaining reconstruction accuracy through the hash encoding of multi-resolution cascades and optimized ray-marching techniques.
@@ II-B Neural Scene Representations
Other methods focus on utilizing explicit or hybrid scene representations for acceleration. Some researchers employ sparse grids for fast inference [37, 71, 59]. Plenoxels [19] extracts a NeRF variant into a sparse voxel grid with spherical harmonics, which enables significantly faster training through direct optimization of explicit representations without neural networks. TensoRF [7] models the radiance field as a 4D tensor and factorizes it into low-rank vectors and matrices, reducing both memory usage and reconstruction time. MINE [31] integrates NeRF with Multiplane Images (MPI). Given a single input image, MINE predicts color and volume density at arbitrary depths and renders novel views using homography warping, enabling efficient reconstruction. Due to the inherent sparsity of 3D scenes, volume representations often suffer from redundancy. To address this, ED3G [5] proposes a compact triplane representation that stores features on axis-aligned planes and employs a lightweight decoder for efficient 3D generation. Similarly, TriMipRF [23] constructs a triplane representation parameterized by three mipmaps, substantially improving training speed while maintaining anti-aliasing effects. However, these triplane representations rely on dense input views. In contrast, our proposed method addresses the challenge under sparse-view constraints and achieves efficient reconstruction. 3DGS [26], a recently developed reconstruction method, achieves high-quality real-time rendering through direct optimization of 3D Gaussian primitives. However, these methods still rely on dense input images for reconstruction and cannot achieve few-shot rendering.
II-C Few-shot Novel View Synthesis
@@
Alternative approaches focus on incorporating regularization constraints [67, 57, 28] or additional supervision [17, 53, 64]. RegNeRF [44] introduces local depth smoothness constraints as regularization to enhance scene structure. DietNeRF [24] employs a pre-trained Vision Transformer (ViT) to enforce semantic consistency across views. FreeNeRF [69] implements frequency regularization to anneal positional encoding, enabling few-shot novel view synthesis with minimal computational overhead. DS-NeRF [17] utilizes sparse point clouds for depth supervision to ensure correct scene geometries. FrugalNeRF [33] selects pseudo ground truth depth based on the cross-scale reprojection errors to guide its multi-scale weight-sharing voxels in efficiently representing the scene details. However, methods based on regularization face challenges in balancing quality and efficiency, and some methods require specific designs.
Recently, 3D Gaussian Splatting has gained significant popularity for its remarkable progress in rendering quality and real-time performance. However, it still suffers from a strong dependency on the number of input views. To address the challenge of sparse-view inputs, methods such as FSGS [77] and DNGaussian [30] introduce monocular depth estimation priors, combined with depth regularization and point cloud densification strategies to enhance the rendering quality. CoR-GS [75] proposes to jointly train two Gaussian radiance fields, leveraging their mutual inconsistency for collaborative regularization. DropGaussian [46] introduces a structural regularization technique that randomly removes subsets of Gaussians during training to improve the visibility and update frequency of the remaining ones. However, in remote sensing scenes, it will be difficult to guarantee the reliability of monocular depth estimation. Moreover, the carefully designed regularization strategies employed during Gaussian optimization considerably increase model complexity, thereby restricting their applicability. In contrast, our proposed method achieves efficient and high-quality novel view synthesis from sparse viewpoints without any additional priors.
II-D Reconstruction of Remote Sensing Imagery
Unlike object-level or common datasets, remote sensing imagery provides a distinctive bird’s-eye perspective. Owing to the motion of platforms, multi-view remote sensing images typically exhibit sparse viewpoints and restricted viewing angles. These inherent properties of remote sensing data introduce new challenges for 3D reconstruction and interpretation tasks.
According to the data acquisition source, reconstruction targets can be broadly categorized into satellite remote sensing imagery and aerial remote sensing imagery. Satellite image reconstruction primarily focuses on accurately restoring Digital Surface Models (DSM) or performing surface reconstruction. By integrating the Rational Polynomial Coefficient (RPC) imaging model, such methods handle variations in illumination, shadowing, surface features, and seasonal changes across multi-temporal observations. Representative approaches include S-NeRF [18], Sat-NeRF [38], and EO-NeRF [39], among others. In contrast, aerial remote sensing imagery, which is typically collected using unmanned aerial vehicles, emphasizes the quality of novel view rendering. ImMPI [65] proposes to reconstruct scenes using a combination of Multi-Plane Image (MPI) representation and implicit neural representation, while large-scale pretraining is employed to obtain 3D scene priors that accelerate the optimization process. It also introduces LEVIR-NVS, a novel-view synthesis dataset for remote sensing. MPNeRF [21] incorporates multi-plane priors to guide NeRF training, exploiting the advantages of MPI to effectively avoid common pitfalls encountered by NeRF in sparse-view aerial scenes. BungeeNeRF [66], on the other hand, adopts a progressive multi-level supervision framework to achieve multi-scale reconstruction of large-scale urban environments. However, these methods either struggle to handle sparse-view inputs or fall short in terms of reconstruction efficiency and rendering quality.
Furthermore, remote sensing scene reconstruction plays an indispensable role in numerous 3D interpretation tasks. For instance, IRT [48] and R3S [49] utilize neural radiance fields and 3D Gaussian Splatting to represent remote sensing scenes, achieving high-precision and view-consistent semantic segmentation across multiple viewpoints. Therefore, we believe that developing an efficient and high-performance 3D representation in remote sensing scenes is of critical importance for both accurate reconstruction and downstream applications.
@@ II-D Reconstruction of Remote Sensing Imagery
Figure 2: Overview of the proposed TriDF. We introduce an efficient hybrid representation for few-shot novel view synthesis, which takes in sparse posed images and performs fast reconstruction. TriDF consists of a TriPlane branch and a Density Fields branch and separately predicts color and volume density. We also integrate the image-based rendering framework and depth-guided optimization based on 3D point clouds for such hybrid representations to stabilize the training process of few-shot NVS tasks.
@@ III-A Overview
Given a set of sparse multi-view aerial remote sensing images with their corresponding camera poses, we aim to efficiently reconstruct the scene for rendering images from arbitrary novel views. An overview of our proposed method is shown in Fig. 2. To achieve efficient scene reconstruction, we propose a hybrid 3D representation that takes advantage of explicit triplane and implicit density fields, which separately predict the color and volume density of sampling points. High-frequency color features are mapped onto the triplane for direct optimization, and volume density is modeled as continuous density fields through MLPs, which significantly improves training efficiency. Although this hybrid representation accelerates training, its performance drops considerably when input views are sparse. To address this limitation, we incorporate the density fields into an image-based rendering framework to aggregate useful information. Furthermore, we implement a depth-guided optimization approach leveraging point clouds to improve the model’s robustness against sparse inputs while ensuring computational efficiency throughout the training process.
@@
A significant limitation of neural radiance fields with sparse input is their tendency to overfit the training views and fail to reconstruct accurate geometries, resulting in blurry synthesized images. To address this limitation, we propose a depth supervision based on point clouds. The approach begins by generating point clouds through Structure-from-Motion (SfM) and Patch Match Stereo (PMS) [55, 3]. Through re-projection, we extract key points and their associated depth information from the generated point clouds. These key points function as spatial anchors during the training process, providing additional geometric constraints. This supervision effectively stabilizes the optimization process since the incorporation of depth priors helps maintain geometric consistency across different viewpoints.
@@ III-B Triplane-Accelerated Density Fields
The triplane is an efficient 3D scene representation. It obtains corresponding feature vectors by projecting 3D points onto orthogonal feature planes and retrieves the color and volume density through MLP after feature aggregation. A key advantage of this representation is that it enables direct optimization of explicit triplane features, which allows for the use of lightweight MLP decoders that significantly accelerate reconstruction. Our research aims to extend the triplane representation to novel view synthesis tasks with sparse input while addressing two primary challenges: 1) Inaccurate modeling of volume density: The discrete nature of triplane representation makes it challenging to model fine scene geometry, particularly when limited supervision is provided. 2) Integration difficulties with the image-based rendering framework: The triplane representation struggles to effectively leverage this prior, leading to overly smoothed synthesis results.
@@ III-B1 Triplane Representation
To process these features, we employ a compact MLP that combines the triplane features ${f}_{\mathrm{tri}}$ and density features ${f}_{\mathrm{m}}$ to generate intermediate features. Subsequently, we encode view directions on the spherical harmonics basis $\mathrm{SH}(\cdot)$. The encoded directions and intermediate features ${f}_{\mathrm{base}}$ are then processed through the Color MLP to predict the colors. This pipeline can be formally expressed as:
@@ III-B2 Density Fields
where $\gamma(\cdot)$ represents positional encoding and ${f}_{\mathrm{m}}$ denotes density features. Due to spatial continuity, the volume density distribution along a ray is relatively simple compared to color, which often contains complex high-frequency components such as intricate textures. Therefore, modeling volume density alone is more efficient than simultaneously predicting both color and density, which leads to faster convergence.
@@
$\mathcal{L}_{\mathrm{color}}=\mathbb{E}_{\mathbf{r}\in\mathcal{R}_{i}}\|\hat{{C}}(\mathbf{r})-{C}(\mathbf{r})\|_{2}^{2},$ (8)
where $\mathcal{R}_{i}$ indicates the set of sampled rays.